Order: Ampicillin 1g IV every 12 hours. Supply: Ampicillin 1g vial. Directions: Dilute with 4.5ml of sterile saline for a total volume of 5ml. Dilute in 50ml of normal saline and infuse over 15 min. How many ml/hour will you set thr IV pump ?
Answer #1

Total amount of solution after dilution of ampicillin is 5 ml which contain 1gm ordered ampicillin drug.

We need to again dilute in 50 ml NS 0.9% then= 50+5

=55ml total solution we have

Infusion time in minutes is 15 minutes

Then =55÷15

=3.67ml per minute we will give

Question asked for ml/hour setting in Iv pump

3.67×60= approx 220 ml/hour we will start the IV pump

If you have any query we will discuss in comment section otherwise please like Thanks